3. Suppose you are interested in estimating the average salary for biostatisticians in Connecticut. You obtain a random sample of 25biostatisticians hired during the last 5 years. The sample data appeared to be approximately normally distributed with no outliers. The average salary for the sample is $45000 and the sample standard deviation is $3150.

A) Construct a 95%confidence interval for µ the true average salary for newly hired biostatisticians in Connecticut.

B) How large of a sample is needed to estimate µ within $2000?
1
Expert's answer
2020-04-17T18:02:32-0400

The provided sample mean is "\bar{X}=45000" and the sample standard deviation is "s=3150."

A) The size of the sample is "n=25" and the required confidence level is 95%.

The number of degrees of freedom are "df=25-1=24," and the significance level is  "\alpha=0.05."

Based on the provided information, the critical t-value for "\alpha=0.05" and "df=24" degrees of freedom is "t_c=2.064."

The 95% confidence for the population mean "\mu" is computed using the following expression


"CI=(\bar{X}-{t_c\cdot s \over \sqrt{n}}, \bar{X}+{t_c\cdot s \over \sqrt{n}})"

Therefore, based on the information provided, the 95 % confidence for the population mean "\mu" is


"CI=(45000-{2.064\cdot 3150 \over \sqrt{25}}, 45000+{2.064\cdot 3150 \over \sqrt{25}})="

"=(43699.68, 46300.32)"

B) The required confidence level is 95%.


"{t_c\cdot s \over \sqrt{n}}\leq2000"

The critical t-value for "\alpha=0.05" and "df=11" degrees of freedom is "t_c=2.200985"

"{t_c\cdot s \over \sqrt{n}}={2.200985\cdot 3150 \over \sqrt{12}}\approx2001.41>2000"

The critical t-value for "\alpha=0.05" and "df=12" degrees of freedom is "tc=2.178813."


"{t_c\cdot s \over \sqrt{n}}={2.178813\cdot 3150 \over \sqrt{13}}\approx1903.53<2000"

"n=13"